# Actinide signatures in low electron fraction kilonova ejecta

### v1

@@ 4.3 Broadband lightcurves
In general, the B16A and B16B models predict several key actinide features that produce significant impacts on the emergent spectra. As previously noted however, the accuracy of these features is for the most part difficult to verify, with the exception of the late time Ac iii emission. Therefore, a complementary approach to the identification of actinide signatures in KNe may be taken by studying the broadband LCs of the models, which are less susceptible to the exact wavelengths of key features, and generally more representative of the overall SED shape and evolution.
Figure 10: Broadband lightcurves of the models in the UV (top panel), optical (middle panel) and IR (bottom panel). The points correspond to the B16A model, the crosses to the B16B model, and the ‘Y’-shaped markers to the W14 model.
[…]
These choices are motivated by the spectral range of our radiative transfer simulations, which have been set to 5 micron in order to include the 3.6 and 4.5 micron bands of IRAC, of relevance to late time observations of AT2017gfo (Kasliwal et al., 2022). While many facilities observed AT2017gfo across other wavelengths, we have chosen the HST for the UV, and GROND for the optical/NIR in order to minimize the number of different instruments considered.
@@
Analysing first the overall trends in each waveband, we find that the actinide-free model of W14 evolves qualitatively differently to the actinide-rich models. This model shows an initial decline in every waveband, followed by a relatively flat evolution, or even a slight increase in apparent magnitude for optical and UV wavebands. This is most noticeable for bluer wavebands, and corresponds to the spectra becoming progressively bluer with time due to rising temperatures and decreasing optical depths in homogeneous composition models (Pognan et al., 2023). The two Spitzer MIR bands however, show a continuous decrease in magnitude.
@@
In contrast, the actinide-rich B16A and B16B models show a relatively different evolution. After an initial decrease, the UV and bluer optical bands tend to show a slight bump, or at least a plateau, between 20 - 40 days. Following this, a steady decrease in magnitude is found across all wavebands, the exact rate varying from band to band. This bump or plateau found in some of the broadband LCs of the actinide-rich models likely arises from flux moving bluewards as time progresses. However, in contrast to the W14 model which prolongs this plateau, or even yields a brightening in bluer wavebands, the actinide-rich models fade in every waveband past the $\sim 40$ day mark.
@@
In both the B16A and B16B cases, this rate of decline is greater than that of the W14 model, such that the latter appears brighter than the B16B model in certain wavebands at the end of the timespan studied here (e.g. z-band from 80 days onwards). The B16A model typically remains brighter in every waveband however, likely due to the overall greater power output of the nuclear network. The generally hotter temperatures from greater radioactive power of the actinide-rich models leads them to be comparatively more blue than the W14 model. This effect is somewhat opposite of that expected by consideration of expansion opacities (e.g. Flörs et al., 2023; Fontes et al., 2023), however it is not fully clear how the red-to-blue evolution will present itself in cases of inhomogeneous composition, 3D models. As such, we present this comparison fully in the context of the models studied here.
@@
The key differences in the evolution of the broadband LCs between the actinide-rich and free models is likely linked not only to the different model compositions, but also to the different temperature evolutions. Looking at Fig. 3, we see that the actinide-rich models start to cool around 40 - 60 days, whereas the W14 model continues to become hotter with time. The combination of this temperature increase with the dropping optical depth at bluer wavelengths allows the actinide-free model to efficiently transfer flux bluewards. In the actinide-rich models however, the dropping temperatures work against this process, such that it is less efficient. This factor, combined with the overall decreasing luminosity as nuclear energy input drops, leads to the steady fading of the B16A/B models across every waveband.
These results indicate that actinide-rich and actinide-free models may be distinguished by the evolving shape of their broadband LCs, not only arising from different compositions, but perhaps mainly from different resultant nuclear power and therefore corresponding temperature evolution.
@@ 5 Discussion and Conclusion
We provide a first attempt at constraining possible actinide signatures in KN spectra across a broad range of time-scales and wavelengths. We find that our actinide atomic data as calculated by fac targeting completeness over accuracy, yields transitions that are for the most part highly inaccurate, or unverifiable due to lack of reference data in the literature. The exceptions to this are two transitions from Ac iii, which we find to be highly accurate, but the features of which remain subdominant in our model. Our U iv transitions are found to be too red, with experimental sources suggesting that this species should be playing a larger role in the absorption and reprocessing of blue photons by fluorescence. By considering our key Th iii transitions at their real wavelengths in the mid-IR, and scaling the A-values for these transitions following the dipole scaling rule, we find that they may be particularly good observational targets for JWST. Further studies with improved atomic data, ejecta models, and covering the entire spectral range of JWST are required in order to determine if these low lying E1 transitions of Th iii are robust observable spectral actinide signatures.
Given the limitations of establishing the exact accuracy of actinide signatures in spectra, we also consider the photometric evolution of low $Y_{e}$ ejecta by examining broadband LCs. Here, we show that for a given electron fraction, actinide-bearing ejecta evolve qualitatively differently to actinide-free ejecta across all wavelengths, from the near UV to the IR. This is found to be importantly linked to the differing temperature evolution of the ejecta, as well as compositional variations.
@@ Acknowledgements
The authors thank S. Sim for useful discussion and comments on this work. The radiative transfer simulations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), and the Swedish National Infrastructure for Computing (SNIC), at the Parallelldatorcentrum (PDC) Center for High Performance Computing, Royal Institute of Technology (KTH), partially funded by the Swedish Research Council through grant agreement no. 2022-06725. MRW acknowledges supports from the National Science and Techonology Council, Taiwan under Grant No. 111-2628-M-001-003-MY4, the Academia Sinica under Project No. AS-CDA-109-M11, and the Physics Division of the National Center for Theoretical Sciences, Taiwan. GMP and AF acknowledge support by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (ERC Advanced Grant KILONOVA No. 885281), the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project-ID 279384907 - SFB 1245, and MA 4248/3-1, and the State of Hesse within the Cluster Project ELEMENTS. AJ acknowledges funding by the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Program (ERC Starting Grant 803189 – SUPERSPEC), the Swedish Research Council (grant 2018-03799), and the Knut and Alice Wallenberg Foundation ("Gravity Meets Light"). RFS acknowledges the support from National funding by FCT (Portugal), through the individual research grant 2022.10009.BD and through project funding 2022.06730.PTDC, "Atomic inputs for kilonovae modeling (ATOMIK)". J.G. acknowledges support from the Swedish Research Council through the individual project grant with contract no. 2020-05467.

### v2

@@ 4.3 Broadband lightcurves
In general, the B16A and B16B models predict several key actinide features that produce significant impacts on the emergent spectra. As previously noted however, the accuracy of these features is for the most part difficult to verify, with the exception of the late time Ac iii emission. Therefore, a complementary approach to the identification of actinide signatures in KNe may be taken by studying the broadband LCs of the models, which are less susceptible to the exact wavelengths of key features, and generally more representative of the overall SED shape and evolution. It should be kept in mind, however, that in the case of highly inaccurate features, such as those for Th iii in the fac data, the usage of photometry still suffers from this underlying inaccuracy. In this case, the magnitude of the bands around 3 micron where Th iii is most prominently emitting, particularly in the B16B model, may be somewhat overestimated. However, this is not expected to overly change the general evolution of the broadband LCs.
We conduct synthetic photometry on the emergent spectral series from near UV wavelengths to $5\mu$m in the IR, the results of which are shown in Fig. 10. The top-left panel shows the near UV bands, for which we have used the Hubble Space Telescope (HST) wide-field camera’s F225W, F275W and F336W filters. The top-right panel shows optical bands, where we have used the g,r,i,z bands of the Gamma-Ray Burst Optical/Near-Infrared Detector (GROND) instrument, mounted on the MPI/ESO 2.2m telescope at La Silla in Chile [^3] https://www.mpe.mpg.de/~jcg/GROND/. Finally, the bottom panels shows IR bands, where the $1-5\mu$m range is covered using the JWST’s NIRCam wide bands. These choices are motivated by the spectral range of our radiative transfer simulations, which have been set to 5 micron in order to cover the 3.6 and 4.5 micron bands of Spitzer’s IRAC, of relevance to late time observations of AT2017gfo (Kasliwal et al., 2022). We have opted to use JWST’s NIRCam for the IR photometry however, as this represents the cutting edge instrument that has and will be used for IR KNe observations in the future.
@@
Analysing first the overall trends in each waveband, we find that the actinide-free model of W14 generally evolves qualitatively differently to the actinide-rich models. This model shows an initial decline in most wavebands, followed by a relatively flat evolution, or even a slight increase in apparent magnitude for optical and UV wavebands. This is most noticeable for bluer wavebands, and corresponds to the spectra becoming progressively bluer with time due to rising temperatures and decreasing optical depths in homogeneous composition models (Pognan et al., 2023). The reddest bands ($\lambda\gtrsim 2\mu$m), however, evolve in qualitatively similar ways across all models, with the actinide-free model having a somewhat slower decline past 40 days.
@@
Aside from these reddest bands, the actinide-rich B16A and B16B models show a relatively different evolution. After an initial decrease, the UV and bluer optical bands tend to show a slight bump, or at least a plateau, between 20 - 40 days. Following this, a steady decrease in magnitude is found across all wavebands, the exact rate varying from band to band. This bump or plateau found in some of the broadband LCs of the actinide-rich models likely arises from flux moving bluewards as time progresses. However, in contrast to the W14 model which prolongs this plateau, or even yields a brightening in bluer wavebands, the actinide-rich models fade in every waveband past the $\sim 40$ day mark.
@@
In both the B16A and B16B cases, this rate of decline is greater than that of the W14 model, such that the latter appears brighter than the B16B model in certain wavebands at the end of the timespan studied here (e.g. JWST bands from $\sim$80 days onwards). The B16A model typically remains brighter in every waveband however, likely due to the overall greater power output of the nuclear network. The generally hotter temperatures from greater radioactive power of the actinide-rich models leads them to be comparatively more blue than the W14 model. This effect is somewhat opposite of that expected by consideration of expansion opacities (e.g. Flörs et al., 2023; Fontes et al., 2023), however it is not fully clear how the red-to-blue evolution will present itself in cases of inhomogeneous composition, 3D models. As such, we present this comparison fully in the context of the models studied here.
@@
The key differences in the evolution of the broadband LCs between the actinide-rich and free models is likely linked not only to the different model compositions, but also to the different temperature evolutions. Looking at Fig. 3, we see that the actinide-rich models start to cool around 40 - 60 days, whereas the W14 model continues to become hotter with time. The combination of this temperature increase with the dropping optical depth at bluer wavelengths allows the actinide-free model to efficiently emit blue photons that can escape the ejecta more easily. In the actinide-rich models however, the dropping temperatures work against this process, such that it is less efficient. This factor, combined with the overall decreasing luminosity as nuclear energy input drops, leads to the steady fading of the B16A/B models across every waveband.
These results indicate that actinide-rich and actinide-free models may be distinguished by the evolving shape of their broadband LCs, not only arising from different compositions, but perhaps mainly from different resultant nuclear power and therefore corresponding temperature evolution. From Figure 10, we suggest that the near UV or optical LCs may provide the best possibility of establishing the significant presence or absence of actinide-bearing ejecta, given that the model LCs behave increasingly similarly the longer the wavelengths considered, such that their evolution is qualitatively similar in wavebands past $\sim 2\mu$m. The caveat here is that the near UV and optical bands are typically fainter than the IR bands, which may present a limitation on the application of this method to distant or relatively faint KNe.
@@ 5 Discussion and Conclusion
We provide a first attempt at constraining possible actinide signatures in KN spectra across a broad range of post-photospheric epochs and wavelengths. We find that our actinide atomic data as calculated by fac targeting completeness over accuracy, yields transitions that are for the most part highly inaccurate, or unverifiable due to lack of reference data in the literature. The exceptions to this are two transitions from Ac iii, which we find to be highly accurate, but the features of which remain subdominant in our model. Our U iv transitions are found to be too red, with experimental sources suggesting that this species should be playing a larger role in the absorption and reprocessing of blue photons by fluorescence. By considering our key Th iii transitions at their real wavelengths in the mid-IR, and scaling the A-values for these transitions following the dipole scaling rule, we find that they may be particularly good observational targets for JWST. Further studies with improved atomic data, ejecta models, and covering the entire spectral range of JWST are required in order to determine if low lying E1 transitions of Th iii are robust observable spectral actinide signatures.
Given the limitations of establishing the exact accuracy of actinide signatures in spectra, we also consider the photometric evolution of low $Y_{e}$ ejecta by examining broadband LCs. Here, we show that for a given electron fraction, actinide-bearing ejecta evolve qualitatively differently to actinide-free ejecta across most wavelengths, from the near UV to the IR up to about 2 micron. This is found to be importantly linked to the differing temperature evolution of the ejecta, as well as compositional variations.
@@ Acknowledgements
The authors thank S. Sim for useful discussion and comments on this work, as well as the anonymous referee for thorough and helpful comments. The radiative transfer simulations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), and the Swedish National Infrastructure for Computing (SNIC), at the Parallelldatorcentrum (PDC) Center for High Performance Computing, Royal Institute of Technology (KTH), partially funded by the Swedish Research Council through grant agreement no. 2022-06725. MRW acknowledges supports from the National Science and Techonology Council, Taiwan under Grant No. 111-2628-M-001-003-MY4, the Academia Sinica under Project No. AS-CDA-109-M11, and the Physics Division of the National Center for Theoretical Sciences, Taiwan. GMP and AF acknowledge support by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (ERC Advanced Grant KILONOVA No. 885281), the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project-ID 279384907 - SFB 1245, and MA 4248/3-1, and the State of Hesse within the Cluster Project ELEMENTS. AJ acknowledges funding by the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Program (ERC Starting Grant 803189 – SUPERSPEC), the Swedish Research Council (grant 2018-03799), and the Knut and Alice Wallenberg Foundation ("Gravity Meets Light"). RFS acknowledges the support from National funding by FCT (Portugal), through the individual research grant 2022.10009.BD and through project funding 2022.06730.PTDC, "Atomic inputs for kilonovae modeling (ATOMIK)". J.G. acknowledges support from the Swedish Research Council through the individual project grant with contract no. 2020-05467.